OpenStudy (anonymous):

Get a medal :3 What is (3.4 x 10^2) x (1.3 x 10^3)? A. 4.42 x 10^5 B. 4.7 x 10^5 C. 4.42 x 10^6 D. 4.7 x 10^6

OpenStudy (akashdeepdeb):

(3.4 x 10^2) x (1.3 x 10^3)? As we have to multiply here. Just multiply the 10^(x) individually and the 3.4 and 1.3 individually. The answer would be.. something like this (3.4 x 10^2) x (1.3 x 10^3) (3.4 * 1.3) * (10^2 * 10^3) (4.42) * (10^(2+3)) = 4.42 * 10^5 [Because \(a^m * a^n = a^{m+n}\)] Understood? :)

OpenStudy (anonymous):

@AkashdeepDeb Yep, perfectly understand. That's the same answer I came up with. Thank you my good sir :3

OpenStudy (akashdeepdeb):

:)